A motor boat traveled 12 miles with the current, turned around, and returned 12 miles against the current to its starting point. The trip with the current took 2 hours and the trip against the current took 3 hours. Find the speed of the boat and the speed of the current.
Answer by checkley77(12844)

DISTANCE=RATE*TIME
T=D/R
2=12/R
2R=12
R=6 MPH. FOR THE 2 HOUR TRIP.
3=12/R
3R=12
R=12/3
R=4 MPH. FOR THE 3 HOUR TRIP.
(6-4)/2=2/2=1 MPH IS THE CURRENT.
SPEED OF THE BOAT IS 6-1=5 MPH. FOR THE BOAT
OR 4+1=5 MPH. FOR THE BOAT